1. Suppose the market for oranges is perfectly competitive and unregulated. Suppose also that the chemicals used to keep the oranges insect-free damage the environment by an estimated $1 per bushel of oranges. Suppose QD = 1000 - 100P and QS = -100 + 100P. Calculate the market equilibrium quantity. (3 marks)


2. Calculate the price consumers would have to pay for the market to achieve the socially optimal level of production. [Hint: use QS = QS + tax but firms will receive a price P – t, therefore, QS = -100 + 100 (P-1). Find new equilibrium price] (3 marks)


3. Calculate the market equilibrium quantity. (1 mark)


4. Calculate the tax revenue collected by the government (1 mark)


1
Expert's answer
2021-06-14T15:22:45-0400

Solution


1. Qd= 1000-100p=Qs=-100+100p


1000-100p=-100+100p

1100=200p

P=5.5

Q=-100+100(5.5)

Q= 450 units


2.Qs=Qs+tax

Qs=-100+100(p-1)

Q=450

450=-100+100(p-1)

450=-100+100p-100

650=100p

P=6.5


3.Qs=-100+100p

P=6.5

-100+100(6.5)

Qs= 550 units


4. Qs=Qs+tax

550=450+tax

Tax=550-450

Tax= 100